Vigilance / Low pressure Board Completed on 21.01.12. - ready for latter wiring.
Vigilance /Low Pressure Board wired on 10.05.12 - Head Lamp switch still to be wired.
Head Lamp cables attached and switch plate refitted on 24.05.12.
Wiring completed on 2nd Man's Driver Controller on 31.03.12.
Completed wiring now includes the foot heater and dead man's pedal on 26.04.12.
Cover Plate for Foot heater and Deadman's Pedal replaced on 04.05.12.
Wiring of Deadmans Pedal and Foot Heater complete on 22.08.12. Only now needs Cover Plate to be fitted.
Start made in refitting Driver's deadman's cover plate on 24.05.12
Cover and Dead man's Pedal replaced on 06.09.12.
More wiring took place on 31.05.12.
Wiring of Driver's Power Controller reaches halfway stage on 23.06.12.
Wiring of Driver's Controller completed on 01.07.12.
Wiring of Wiring of Centre Console has commenced on 06.11.12.
Wiring of Console on 08.12.12 having continued in past month.
Wiring of Centre Console continued on 08.12.12. Just 3 more switchs to do.